An emf is induced in a conducting loop of wire 1.45 m long as its shape is changed from square to circular. Find the average magnitude of the induced emf if the change in shape occurs in 0.125 s and the local 0.479-T magnetic field is perpendicular to the plane of the loop.
Answer #1

Faraday's Law: N=1
Emf = Δflux/Δt = -N*B*ΔA/Δt
The area of the loop changes
from a square to a circle in 0.125 s

B = 0.479 T

L = 1.45 m
r = L / (2 * π) m
A_c = π * r ^ 2 m^2 (area circle)
A_s = (L / 4) ^ 2 m^2 (area square)
Δt = 0.125 s

Emf = -N*B*ΔA/Δt
Emf = -B * (A_s - A_c) / Δt
Emf= -(0.479)*(0.13 - 0.16)/0.125

Emf= 0.115 volts